OpenStudy (dboyette):

in testing a new drug, researchers found that 10% of all patients using it will have a mild side effect. A random sample of 12 patients using the drug is selected. Find the probability that A. exactly three will have this mild side effect B. at least three will have this mild side effect

OpenStudy (amistre64):

is 12 enough of a count?

OpenStudy (amistre64):

(12 C 3) .10^3 .90^9 maybe?

OpenStudy (amistre64):

at least 3 is 1-"2"

OpenStudy (amistre64):

if i hit the right buttons, id say A is .0852

OpenStudy (amistre64):

B might be .7699 if im remembering it right

OpenStudy (amistre64):

there he is :)

OpenStudy (anonymous):

binomial this one

OpenStudy (amistre64):

yay!! i was right, basically

OpenStudy (dboyette):

exactly three is 0.85 and at least 3 would be 0.025? Am i right?

OpenStudy (anonymous):

\[P(X=3)=\dbinom{12}{3}(.1)^3(.9)^9\]

OpenStudy (amistre64):

1 - exactly2

OpenStudy (amistre64):

.... or am i not hitting the right synapses

OpenStudy (amistre64):

1-P(1)-P(2) is what i got in me head

OpenStudy (anonymous):

at least three compute none \[(.9)^{12}\] 1 \[12\times .1\times (.9)^{11}\] and 2 \[66(.1)^2(.9)^{10}\] add them up and subtract the result from 1

OpenStudy (anonymous):

@ amistre you forgot \[P(X=0)\] you need that one too

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

\[P(X\geq 3)=1-(P(X=0)+P(X=1)+P(X=2))\]

OpenStudy (anonymous):

let me get a number for the first one

OpenStudy (anonymous):

.085 not .85

OpenStudy (dboyette):

yes i see that now